OpenStudy (anonymous):

find all zeros of the polynomial function f(x)=x^4-6x^3+7x^2+6x-8

OpenStudy (anonymous):

i am not completely clueless i have an idea of it but im stuck on a step:S

OpenStudy (campbell_st):

the factors of the constant term are call them p \[\pm 1, \pm 2, \pm 4, \pm 8\] the factors of the leading term are call them q \[\pm 1\] so the rational factors are p/q so try x = -1 \[f(-1) = (-1)^4 - 6(-1)^3 + 7(-1)^2 + 6(-1) - 8\] f(-1) = 0 so then (x + 1) now try x = 1 f(1) = 0 so ( x - 1) is a factor this means x^2 -1 is also a factor so do some polynomial division  this should result in a quadratic factor.. which could then be factorised to find the linear factors and hence all the zeros.